OpenStudy (dan815):

Can you write a general expression for the derivative of the product of n functions

OpenStudy (dan815):

@Empty

OpenStudy (anonymous):

keep moving the primes along

zepdrix (zepdrix):

Hmm I know you can write the nth derivative in a fancy way. it follows the same type of format as the binomial theorem. But n functions? Hmm interesting idea :)

OpenStudy (dan815):

lol did u know i wanted this because of arithmetic derivatives =.=

OpenStudy (empty):

They're direct messages, not private messages :O

OpenStudy (empty):

dan's being watched (and for good reason I might add)

OpenStudy (empty):

ohhhh satellite is talking about like \[(fgh)'=f'gh+fg'h+fgh'\] moving these primes along dan haha

OpenStudy (dan815):

every permutations of the function - 1 function must exist i see

OpenStudy (empty):

\[(abc)'=(ab)'c+abc'=(a'b+ab')c+abc'=a'bc+ab'c+abc'\] Actually the derivative for the arithmetic functions will have the same general form I think FOR THOSE WHO AREN'T ME OR DAN LISTEN UP! We're playing around with the arithmetic derivative itself found here: https://projecteuler.net/problem=484 (We're not attempting the problem though, maybe another day haha)

OpenStudy (anonymous):

oops

OpenStudy (empty):

Feel free to join in it's super fun to play with lol. dan, here's my hint to you, divide by fgh and you get this: \[\frac{(fgh)'}{fgh} = \frac{f'}{f}+\frac{g'}{g} + \frac{h'}{h}\]

OpenStudy (dan815):

hey i think i got it how about hte greatest GCD is all the primes -1 exponent

OpenStudy (empty):

\[\gcd(4,4')=?\]

OpenStudy (dan815):

(4,(2*2)')=2

OpenStudy (empty):

do it again slower

OpenStudy (dan815):

okay wait maybe we need another rule

OpenStudy (dan815):

it should be 4 so this isnt working lol

OpenStudy (dan815):

okay how about only for p1^p1 that is different

OpenStudy (empty):

so you know the power rule, \[(p^k)' = kp^{k-1}\] what numbers solve this differential equation: \[n'=n\]

OpenStudy (dan815):

ah so all the ones in form pn^pn

OpenStudy (dan815):

when k=p

OpenStudy (empty):

Yeah so it's pretty cool we have like: \[(p^{kp})' = k p^{kp}\] kinda serving like \[(e^{kx})' =ke^{kx}\] haha Also here's the chain rule: \[(a^k)' = ka^{k-1} a'\] one fun thing to prove; \[1'=?\]

OpenStudy (empty):

\[1=(1*1)\] \[1'=1'*1+1*1'\] \[1'=2*1'\] \[1 \ne 2\] \[\implies\] \[1'=0\]